OpenStudy (anonymous):

i know i already posted this question but i am confused Write the explicit formula that represents the geometric sequence -2, 8, -32, 128

OpenStudy (anonymous):

Explicit formula means that it you plug in a value for x and you immediately get the answer. Therefore, we know the first term is -2, the second is 8, and so on. There is a common ratio of -4. We multiply each term by -4. Therefore, y=-4x + c. We don't know what c is yet. To find c, we plug in the first term. -2=-4(1) + c. This simplifies to 2=c so c is 2. We then add that back to the original equation. y=-4x + 2. This should be your equation

OpenStudy (anonymous):

this is wat furuz told me and then he said it was wrong so i need the right way

geerky42 (geerky42):

Explicit formula for geometric sequence looks like \(a_n = a_1\cdot r^{n-1}\), right?

OpenStudy (anonymous):

yes

geerky42 (geerky42):

And you already knew that \(a_1 = -2\) and \(r=-4\), therefore explicit formula is \(\boxed{a_n = -2(-4)^{n-1}}\)

OpenStudy (anonymous):

correct

geerky42 (geerky42):

Well, there you go.

OpenStudy (anonymous):

so thas my aswer an = -2(-4)^n1

geerky42 (geerky42):

Yeah

OpenStudy (anonymous):

thanxs

geerky42 (geerky42):

No problem